John's paycheck is for $x and y cents. When he cashes it the cashier accidentally gives him $y and x cents. As a result John ( who is not paying attention at all) has lost $95.04. How much was John's original check for. Find all the possible solutions.
Answer by stanbon(75887)

John's paycheck is for $x and y cents. When he cashes it the cashier accidentally gives him $y and x cents. As a result John ( who is not paying attention at all) has lost $95.04. How much was John's original check for. Find all the possible solutions.
-----------------
Equation:
check value - cashier pay = $95.04
100x+y - (100y+x) = 9504 cents
-----------------------
99x-99y = 9504
x - y = $96
---
y = x-96
---
y cannot be negative
x >= 96
96 <= x <= 99
------------------
so x = 96 or 97 or 98 or 99
---
Johns paycheck is $x and y cents =
$96.00
or 97.01
or 98.02
or 99.03
=================
Cheers,
Stan H.
=================